OpenStudy (anonymous):

what are the coordinates of the terminal point determined by θ=7π/6?

OpenStudy (anonymous):

\[- \frac{ \sqrt{3} }{ 2 } , - \frac{ 1 }{ 2 }\] it is on the unit circle

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes those two are the coordinates , look at the unit circle

OpenStudy (anonymous):

thanks :) i'll check it out now

OpenStudy (anonymous):

but - sqrt(3)/2 is not a point, is it?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

It's equal to -.866

OpenStudy (anonymous):

as long as it is in standard position those will be the coordinates

OpenStudy (anonymous):

oh okay, thanks. sorry about that

OpenStudy (anonymous):

its ok n.p